# across languages

### v1

@@ Abstract
As large language models (LLMs) are deployed globally, it is crucial that their responses are calibrated across languages to accurately convey uncertainty and limitations. Previous work has shown that LLMs are linguistically overconfident in English, leading users to overrely on confident generations. However, the usage and interpretation of epistemic markers (e.g., ‘It’s definitely’, ‘I think’) can differ sharply across languages. Here, we study the risks of multilingual linguistic (mis)calibration, overconfidence, and overreliance across five languages to evaluate the safety of LLMs in a global context.
We find that overreliance risks are high across all languages. We first analyze the distribution of LLM-generated epistemic markers, and observe that while LLMs are cross-linguistically overconfident, they are also sensitive to documented linguistic variation. For example, models generate the most markers of uncertainty in Japanese and the most markers of certainty in German and Mandarin. We then measure human reliance rates across languages, finding that while users strongly rely on confident LLM generations in all languages, reliance behaviors differ cross-linguistically: for example, users rely significantly more on expressions of uncertainty in Japanese than in English. Taken together, these results indicate high risk of reliance on overconfident model generations across languages. Our findings highlight the challenges of multilingual linguistic calibration and stress the importance of culturally and linguistically contextualized model safety evaluations.
@@ 1 Introduction
Past work has found that English language models are systematically overconfident—that is, they generate epistemic markers expressing high certainty even when incorrect, with high frequency. Compounding this, English speakers tend to systematically overrely on language model outputs (Zhou et al., 2024b). Yet, in the linguistics literature, it is well-established that the (1) usage and (2) interpretation of epistemic markers differ sharply and richly across languages (Davidson & Chrisman, 1994; Itani, 1995; Doupnik & Richter, 2003; Yang, 2013). Here, we contend that the current literature’s focus on English epistemic markers is insufficient for understanding the unique safety harms multilingual models pose to a global audience. Our contributions are as follows:
@@
    Model overconfidence differs across languages. Although it is well-documented that LLMs generate linguistically confident but incorrect answers in English, it is unclear whether this effect holds across languages. Examining five languages (English, French, German, Japanese, and Mandarin), we find that LLMs are sensitive to distributional differences in the usage of epistemic markers between languages (e.g. more hedges in Japanese than in English), but that overconfidence rates are high across all five languages.
@@
    Humans reliance differs across languages. Thus far, reliance have focused on English-speaking users (Zhou et al., 2024c). Yet in reality, reliance on epistemic markers will vary based on linguistic and cultural norms. In a bilingual human reliance study, we show that, while high in all languages, reliance rates do differ significantly between languages.
@@
    Yet overreliance risks remains high across languages. Given our first result, we might naïvely expect that overreliance risks—i.e. relying on overconfident generations—also differ across languages. For example, since LLMs produce more hedges in Japanese than in English, Japanese speakers might actually be less prone to overreliance risk. But taking into account differences in reliance rates, we find that this reasoning is incorrect. Instead, the same bilingual speaker will systematically rely more on Japanese uncertainty markers than their translated English counterparts. In other words, the distribution of human reliance differs across languages in a way that cancels out the effect of distribution shift in LLM overconfidence.
@@
Taken together, our results implicate significant global safety risks for overreliance. We argue that the appropriate use of epistemic markers necessitates contextualization of expressions to the language of use. While our work here concentrates only on difference in language, we expect that similar contextualization is necessary for differences in topic and setting, as well as other differences in user demographics (e.g. age, gender). Our results highlight the existing risks of LLM overconfidence and the linguistic and social norms that must be considered when developing safe and calibrated language models.
@@ 2.1 LLMs and Uncertainty Expressions
Recent works in NLP have been interested in aligning internal model probabilities with their task accuracy (Jiang et al., 2021; Desai & Durrett, 2020; Jagannatha & Yu, 2020; Kamath et al., 2020; Kong et al., 2020; Hofweber et al., 2024). Most recently, there has been growing interest in calibrating models in a manner more accessible to end users by directly emitting model probabilities as a proxy for model confidence to the user (Kadavath et al., 2022; Tian et al., 2023; Liu et al., 2023; Xiong et al., 2024; Tanneru et al., 2023; Mielke et al., 2022; Lin et al., 2022; Stengel-Eskin et al., 2024). A key finding of this line of work is that English LLMs are miscalibrated to their accuracies: models are less accurate when they produce markers of certainty. Zhou et al. (2023) posit that this is an artifact of (1) pretraining data, which contains a large number of Q/A pairs from online forums in which questions are highly confident and answers often include hedges, and (2) the preference optimization process, as human users ‘prefer’ more certain responses.
@@ 2.2 Cross-Linguistic Variation
Previous work has found that while English and French speakers tend to use mostly moderate expressions of certainty—‘I’m pretty sure,’ ‘I believe’—speakers of Japanese tend to use weakeners—‘I think,’ ‘maybe’—as a baseline (Davidson & Chrisman, 1994; Itani, 1995; Lauwereyns, 2000; Itakura, 2013; Barotto, 2018; Yin et al., 2024). On the other hand, speakers of German and Mandarin tend to use distributions tending on the side of strengtheners (Doupnik & Richter, 2003; Kranich, 2011; Yang, 2013). These effects are robust across speaker identity and domain. [^1] Similar results have been shown for other languages, including Hindi, Turkish, and Italian (Uysal, 2014; Mazzuca et al., 2024). Here, we concentrate on English, French, German, Japanese, and Mandarin as these distributional effects are well-documented and robust, and these languages are well-supported by multilingual LLMs.
Taking our current understanding of LLM overconfidence, this work seems to suggest new miscalibration risks for non-English languages. For one, it may be the case that overconfidence rates differ between languages—for example, if LLMs produce more strengtheners in German or Mandarin, then perhaps they are more overconfident in these languages (and vice versa in Japanese). Further, cross-linguistic distributional differences may lead to differences in reliance patterns between languages. For example, since weakeners are a priori more frequent in Japanese than in English, a user might correspondingly rely more heavily on a weakener in Japanese. In other words, their ‘internal model’ of the LLM might change between languages (in the sense of Gricean recursive reasoning; see Frank & Goodman, 2012; Goodman & Frank, 2016; Degen, 2023).
@@ 3 LLM Generations Are Overconfident Across Languages
Past work has found that English LLMs are prone to risks stemming from overconfidence—that is, expressing epistemic certainty even when providing incorrect knowledge. In this section, we investigate the global risks of LLMs by assessing their rate of overconfidence in multiple different languages. We show that multilingual LLMs demonstrate the same overconfidence effects in languages other than English and that LLMs are sensitive to variation in the linguistic norms of using uncertainty in different languages.
@@ A Note On LLM Epistemic States
Previous work has suggested that LLMs rarely produce verbalized expressions of confidence, unless explicitly prompted to do so (e.g. Zhou et al., 2024c). Yet, humans tend to rely on ‘plain’ model generations (i.e. those not containing epistemic markers) even more than they do on strengtheners (Zhou et al., 2024b, also see Figure 5). In other words, unprompted models lead to high risks of overreliance. But this can be mitigated: users tend to be sensitive to linguistic markers of uncertainty (i.e. they rely less on a model when it hedges). As such, here we evaluate models by explicitly eliciting epistemic markers as part of their responses using few-shot prompting.
@@ 3.1 Methods
As very few models expressly support multilingual use, we were limited in the breadth of models that could be studied. We analyzed three publicly deployed LLMs at varying scales: GPT-4o (May 2024; Hurst et al., 2024) and Llama 3.1 8B/70B Instruct (July 2024; Grattafiori et al., 2024). [^2] GPT-4o responses were collected in September 2024 for English, German, French, and Japanese, and February 2025 for Mandarin. Note that while all three of these models are designed to be multilingual, the Llama models do not explicitly support Japanese or Mandarin (they do, however, support French and German). As a high-level overview, we first prompted each model to answer multiple-choice questions and to include an epistemic marker in their response. We then annotated these responses by their expressed certainty.
@@
Model English French German Japanese Mandarin
GPT-4o 80.96 76.32 77.83 77.08 75.75
Llama-3.1-70B 66.69 57.51 34.47 45.90 55.65
Llama-3.1-8B 59.44 37.17 23.05 31.29 42.17
@@ 3.3 LLM Generations Adhere to Linguistic Norms
We are also interested in the baseline distribution of generations across languages—in particular, whether this matches documented linguistic norms. We find that models are generally sensitive to cross-linguistic variation in usage of uncertainty expressions (Figure 2). For example, while in English and French we find that most generations include moderate or strong epistemic markers (96% and 91% in GPT-4o, respectively), we see that in Japanese, most epistemic markers are weakeners (59% in GPT-4o). On the other hand, in German and Mandarin, strengtheners are more prevalent than weak or moderate markers (53% and 80% in GPT-4o, compared to 42% in English).
@@ 4 Humans Overrely on LM Generations Across Languages
We find that users overrely on generations including strengtheners in all languages—that is, they systematically rely on strengtheners at a rate leading to high risk of relying on incorrect generations. Further we also find cross-linguistic variation in the distribution of reliance.
@@ 4.1 Methods
The task used 60 trivia questions from Zhou et al. (2024c), which consisted of the 60 most difficult geography questions on Sporcle, an online trivia platform. Using difficult questions encourages users to rely on the model response rather than their own knowledge. We then machine-translated these questions into each target language with the Google Translate API.
@@ 4.2 Overreliance Across Languages
Using this new metric, we find that overreliance risk is high across languages in models. Even in GPT-4o, the highest accuracy model evaluated, we find an average overreliance risk of nearly 10% across languages. The overreliance risk is substantially worse in Llama-3.1-8B, where users are predicted to rely on incorrect responses 40% of the time.
@@ 4.3 Reliance Differences Between Languages
Given that overreliance appears across languages, we might naively assume that, as LLMs produce more hedges in Japanese, Japanese users might face lower risks of overreliance. Yet, we find that this is incorrect: human reliance rates actually differ by category across languages.
@@
Yet in Japanese, we observe that when the model uses any form of epistemic marker, human reliance increases in comparison to English. This leads to an increased risk of overreliance, as even though models produce more hedges in Japanese, speakers of Japanese are still likely to rely on hedges moreso than in English. In other words, the shift in the distribution of epistemic markers does not actually mitigate overreliance risks. We find that this manifests in the overreliance risks in Table 9: Japanese users are 1.5x more susceptible to overreliance risk than English users.
Figure 3: Differences in human reliance between English and each target language, by condition. Error bars represent bootstrapped 95% CI. Reliance rates on plain expressions is lower in German and French than in English. In German, bilingual users rely less on strengtheners than they do in English. Contrastingly, in Japanese, users rely more on all epistemic markers. French epistemic markers are perceived roughly as they are in English.
@@ 5 Discussion and Conclusion
Here, we studied the risks arising from miscalibrated multilingual language models. We found that, while multilingual LLMs do adhere to documented linguistic norms around the production of epistemic markers, they are still systematically overconfident across languages. Further, we showed that users tend to overrely on LLMs in all languages, and that overreliance risks may actually be worse in languages like Japanese, where expressions uncertainty are more common but have diminished function as markers of epistemic state.
More Uncertainty Generations, Less Perceived Uncertainty Our findings illustrate that although languages might differ in their generations of epistemic markers, this does not guarantee a reduction in LLM overreliance. For example, in Japanese generations we see a simultaneous increase in uncertainty markers along with a decrease in the perceived function of hedges—in other words, even though models generate more hedges in Japanese, users are simultaneously more likely to rely on generations containing them. Applying our understanding of how English-speaking users might rely on English markers would have provided us with an inaccurate estimation of overreliance risks of Japanese generations. Our findings show that reliance on the epistemic markers is dependent on the linguistic norms and emphasizes the need to evaluate overreliance in context.
@@ Appendix A Annotation Details
For each language, we first ran a qualifying study on Prolific in which we asked 15 participants—who self-reported C1/C2 proficiency—to annotate 15 model generations on a 1-7 Likert scale of their certainty. As a baseline, we qualitatively coded these 15 model generations according to our schema before running the qualifying study; for each language, we then selected the 6 participants who were most accurate against this baseline to annotate the entire set of generations. From these 6 participants, we used the top 3 annotators by time taken to complete the study. This step allowed us to filter out participants who did not satisfactorily annotate responses. In Mandarin, we included an additional handwriting check, in which participants were asked to write out a simple sentence in Chinese characters (this was an additional filter against bot submissions, which were especially prevalent for Mandarin as it has a small Prolific user base).
@@ Appendix B Reliance Artifacts
Figure 4: An example of a question from the reliance study, using Zhou et al. (2024b)’s Rel-A.I. framework. Users are shown individual geography questions, along with the beginning of a model’s response, containing an epistemic marker. They are then asked whether or not they would rely on the model response.

### v2

@@ Abstract
As large language models (LLMs) are deployed globally, it is crucial that their responses are calibrated across languages to accurately convey uncertainty and limitations. Prior work shows that LLMs are linguistically overconfident in English, leading users to overrely on confident generations. However, the usage and interpretation of epistemic markers (e.g., ‘I think it’s’) differs sharply across languages. Here, we study the risks of multilingual linguistic (mis)calibration, overconfidence, and overreliance across five languages to evaluate LLM safety in a global context. Our work finds that overreliance risks are high across languages.
We first analyze the distribution of LLM-generated epistemic markers and observe that LLMs are overconfident across languages, frequently generating strengtheners even as part of incorrect responses. Model generations are, however, sensitive to documented cross-linguistic variation in usage: for example, models generate the most markers of uncertainty in Japanese and the most markers of certainty in German and Mandarin. Next, we measure human reliance rates across languages, finding that reliance behaviors differ cross-linguistically: for example, participants are significantly more likely to discount expressions of uncertainty in Japanese than in English (i.e., ignore their ‘hedging’ function and rely on generations that contain them). Taken together, these results indicate a high risk of reliance on overconfident model generations across languages. Our findings highlight the challenges of multilingual linguistic calibration and stress the importance of culturally and linguistically contextualized model safety evaluations.
@@ 1 Introduction
Past work has found that English language models are systematically overconfident—that is, they generate epistemic markers expressing high certainty even when incorrect, with high frequency. Compounding this, English speakers tend to systematically overrely on language model outputs (Zhou et al., 2024b). Yet, in the linguistics literature, it is well-established that the (1) usage and (2) interpretation of epistemic markers differ sharply and richly across languages (Davidson & Chrisman, 1994; Itani, 1995; Doupnik & Richter, 2003; Yang, 2013). Here, we contend that the current literature’s focus on English epistemic markers is insufficient for understanding the unique safety harms multilingual models pose to a global audience.
Our contributions are as follows:
@@
    Models are overconfident across languages. It is well-documented that LLMs generate linguistically confident but incorrect answers in English. We show that this effect replicates across four new languages (French, German, Japanese, and Mandarin), i.e. that models frequently generate strengtheners even when incorrect.
@@
    But model confidence still differs across languages. Given that the usage of epistemic markers differs across languages, we also examine whether LLMs are sensitive to these distributional differences. In all five languages, we find that models indeed differ in their overconfidence across languages: for example, they produce more hedges in Japanese than in English.
@@
    Human reliance differs across languages. Thus far, reliance studies have focused on English-speaking participants (Zhou et al., 2024c). Yet we expect reliance on epistemic markers to vary based on linguistic and cultural norms. In a bilingual human reliance study, we show that, while high in all languages, reliance rates do differ significantly between languages.
  4. 4.
    Yet overreliance risks remains high across languages. Given our first result, we might naïvely expect that overreliance risks—i.e. relying on overconfident generations—also differ across languages. For example, since LLMs produce more hedges in Japanese than in English, Japanese speakers might actually be less prone to overreliance risk. But taking into account differences in reliance rates, we find that this reasoning is incorrect. Instead, the same bilingual speakers are systematically more likely to discount Japanese uncertainty markers than their translated English counterparts. In other words, the distribution of human reliance differs across languages in a way that cancels out the effect of distribution shift in LLM overconfidence.
@@
Taken together, our results imply significant global safety risks for overreliance. We argue that the appropriate use of epistemic markers must be contextualized to the language being used. While our work here concentrates only on difference in language, we expect that similar contextualization is necessary for differences in topic and setting, as well as other differences in participant demographics (e.g. age, gender). Our results highlight the existing risks of LLM overconfidence and the linguistic and social norms that must be considered when developing safe and calibrated language models.
@@ 2.1 LLMs and Uncertainty Expressions
Recent work in NLP has focused on aligning internal model probabilities with their task accuracy (Jiang et al., 2021; Desai & Durrett, 2020; Jagannatha & Yu, 2020; Kamath et al., 2020; Kong et al., 2020; Hofweber et al., 2024). Most recently, there has been growing interest in calibrating models in a manner more accessible to end users by directly emitting model probabilities as a proxy for model confidence to the user (Kadavath et al., 2022; Tian et al., 2023; Liu et al., 2023; Xiong et al., 2024; Tanneru et al., 2023; Mielke et al., 2022; Lin et al., 2022; Stengel-Eskin et al., 2024). A key finding of this line of work is that English LLMs are miscalibrated to their accuracies: models are less accurate when they produce markers of certainty. Zhou et al. (2023) posit that this is an artifact of (1) pretraining data, which contains a large number of Q/A pairs from online forums in which questions are highly confident and answers often include hedges, and (2) the preference optimization process, as human users ‘prefer’ more certain responses.
@@ 2.2 Cross-Linguistic Variation
Previous work has found that while English and French speakers tend to use mostly moderate expressions of certainty—‘I’m pretty sure,’ ‘I believe’—speakers of Japanese tend to use weakeners—‘I think,’ ‘maybe’—as a baseline (Davidson & Chrisman, 1994; Itani, 1995; Lauwereyns, 2000; Itakura, 2013; Barotto, 2018; Yin et al., 2024). On the other hand, speakers of German and Mandarin tend to use distributions with morestrengtheners (Doupnik & Richter, 2003; Kranich, 2011; Yang, 2013). These effects are robust across speaker identity and domain. [^1] Similar results have been shown for other languages, including Hindi, Turkish, and Italian (Uysal, 2014; Mazzuca et al., 2024). Here, we concentrate on English, French, German, Japanese, and Mandarin as these distributional effects are well-documented and robust, and these languages are well-supported by multilingual LLMs.
This work offers a new understanding of LLM overconfidence and suggests additional safety risks for overreliance for non-English languages. For one, it may be the case that overconfidence rates differ between languages—for example, if LLMs produce more strengtheners in German or Mandarin, then perhaps they are more overconfident in these languages (and vice versa in Japanese). Further, cross-linguistic distributional differences may lead to differences in reliance patterns between languages. For example, since weakeners are a priori more frequent in Japanese than in English, a participant might correspondingly be more likely to discount a weakener in Japanese. In other words, their ‘internal model’ of the LLM might change between languages (in the sense of Gricean recursive reasoning; see Frank & Goodman, 2012; Goodman & Frank, 2016; Degen, 2023).
@@ 3 LLM Generations Are Overconfident Across Languages
Past work has found that English LLMs are prone to risks stemming from overconfidence—that is, expressing epistemic certainty even when providing incorrect knowledge. In this section, we investigate the global risks of LLMs by assessing their rate of overconfidence in multiple different languages. We show that multilingual LLMs demonstrate the same overconfidence effects in languages other than English and also that LLMs are sensitive to variation in the linguistic norms of using uncertainty in different languages.
@@ A Note On LLM Epistemic States
Previous work has suggested that LLMs rarely produce verbalized expressions of confidence, unless explicitly prompted to do so (e.g. Zhou et al., 2024c). Yet, humans tend to rely on ‘plain’ model generations (i.e. those not containing epistemic markers) even more than they do on strengtheners (Zhou et al., 2024b, also see Figure 5). In other words, unprompted models lead to high risks of overreliance. But this can be mitigated: participants tend to be sensitive to linguistic markers of uncertainty (i.e. they rely less on a model when it hedges). As such, here we evaluate models by explicitly eliciting epistemic markers as part of their responses using few-shot prompting.
Model English French German Japanese Mandarin
GPT-4o 80.96 76.32 77.83 77.08 75.75
Llama-3.1-70B 66.69 57.51 34.47 45.90 55.65
Llama-3.1-8B 59.44 37.17 23.05 31.29 42.17
@@ 3.1 Methods
As very few models expressly support multilingual use, we were limited in the breadth of models that could be studied. We analyzed three publicly deployed LLMs at varying scales: GPT-4o (May 2024; Hurst et al., 2024) and Llama 3.1 8B/70B Instruct (July 2024; Grattafiori et al., 2024). [^2] GPT-4o responses were collected in September 2024 for English, German, French, and Japanese, and February 2025 for Mandarin. Note that while all three of these models are designed to be multilingual, the Llama models do not explicitly support Japanese or Mandarin (they do, however, support French and German). As we describe below, we first prompted each model to answer multiple-choice questions and to include an epistemic marker in their response, and then annotated these responses by their expressed certainty.
@@ 3.3 LLM Generations Adhere to Linguistic Norms
We are also interested in the baseline distribution of generations across languages—in particular, whether model outputs match documented linguistic norms. We find that models are generally consistent with cross-linguistic variation in usage of uncertainty expressions (Figure 2). For example, while in English and French we find that most generations include moderate or strong epistemic markers (96% and 91% in GPT-4o, respectively), we see that in Japanese, most epistemic markers are weakeners (59% in GPT-4o). On the other hand, in German and Mandarin, strengtheners are more prevalent than weak or moderate markers (53% and 80% in GPT-4o, compared to 42% in English).
@@ 4 Humans Overrely on LM Generations Across Languages
We find that participants overrely on generations including strengtheners in all languages—that is, they systematically rely on strengtheners at a rate leading to high risk of relying on incorrect generations. Further we also find cross-linguistic variation in the distribution of reliance.
@@ 4.1 Methods
The task used 60 trivia questions from Zhou et al. (2024c), which consisted of the 60 most difficult geography questions on Sporcle, an online trivia platform. Using difficult questions encourages participants to rely on the model response rather than their own knowledge. We then machine-translated these questions into each target language with the Google Translate API.
@@ 4.2 Overreliance Across Languages
Using this new metric, we find that overreliance risk is high across languages in models. Even in GPT-4o, the highest accuracy model evaluated, we find an average overreliance risk of nearly 10% across languages. The overreliance risk is substantially worse in Llama-3.1-8B, where participants are predicted to rely on incorrect responses 40% of the time.
@@ 4.3 Reliance Differences Between Languages
Given that overreliance appears across languages, we might naively assume that, as LLMs produce more hedges in Japanese, Japanese participants might face lower risks of overreliance. Yet, we find that this is incorrect: human reliance rates actually differ by category across languages.
@@
Yet in Japanese, we observe that when the model uses any form of epistemic marker, human reliance increases in comparison to English. This leads to an increased risk of overreliance, as even though models produce more hedges in Japanese, speakers of Japanese are still more likely to discount hedges than in English. In other words, the shift in the distribution of epistemic markers does not actually mitigate overreliance risks. We find that this manifests in the overreliance risks in Table 9: Japanese participants are 1.5x more susceptible to overreliance risk than English participants.
Figure 3: Differences in human reliance between English and each target language, by condition. Error bars represent bootstrapped 95% CI. Reliance rates on plain expressions is lower in German and French than in English. In German, bilingual participants rely less on strengtheners than they do in English. Contrastingly, in Japanese, participants rely more on all epistemic markers. French epistemic markers are perceived roughly as they are in English.
@@ 5 Discussion and Conclusion
Here, we studied the risks arising from miscalibrated multilingual language models. We found that, while multilingual LLMs do adhere to documented linguistic norms around the production of epistemic markers, they are still systematically overconfident across languages. Further, we showed that participants tend to overrely on LLMs in all languages, and that overreliance risks may actually be worse in languages like Japanese, where expressions of uncertainty are more common but have diminished function as markers of epistemic state.
More Uncertainty Generations, Less Perceived Uncertainty Our findings illustrate that although languages might differ in their generations of epistemic markers, this does not guarantee a reduction in LLM overreliance. For example, in Japanese generations we see a simultaneous increase in uncertainty markers along with a decrease in the perceived function of hedges—in other words, even though models generate more hedges in Japanese, participants are simultaneously more likely to rely on generations containing them. Applying our understanding of how English-speaking participants might rely on English markers would have provided us with an inaccurate estimation of overreliance risks of Japanese generations. Our findings show that reliance on the epistemic markers is dependent on the norms of the language being spoken, and emphasizes the need to evaluate overreliance in context.
@@ Appendix A Annotation Details
For each language, we first ran a qualifying study on Prolific in which we asked 15 participants—who self-reported C1/C2 proficiency—to annotate 15 model generations on a 1-7 Likert scale of their certainty. As a baseline, we qualitatively coded these 15 model generations according to our schema before running the qualifying study; for each language, we then selected the 6 participants who were most accurate against this baseline to annotate the entire set of generations. From these 6 participants, we used the top 3 annotators by time taken to complete the study. This step allowed us to filter out participants who did not satisfactorily annotate responses. In Mandarin, we included an additional handwriting check, in which participants were asked to write out a simple sentence in Chinese characters (this was an additional filter against bot submissions, which were especially prevalent for Mandarin as it has a small Prolific participant base).
@@ Appendix B Reliance Artifacts
Figure 4: An example of a question from the reliance study, using Zhou et al. (2024b)’s Rel-A.I. framework. Participants are shown individual geography questions, along with the beginning of a model’s response, containing an epistemic marker. They are then asked whether or not they would rely on the model response.